OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day?

OpenStudy (anonymous):

x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution ( i think its no solution )

OpenStudy (welshfella):

3|x − 8| + 15 = 18 3|x − 8| = 18 - 15 = 3 |x − 8| = 3/3 = 1 this means that x-8 = 1 x = 9 and x - 8 = -1 x = 7 minimum temperature = 7 and maximum temperature = 9

OpenStudy (anonymous):

thank u can u help me with 2 more? @welshfella

OpenStudy (welshfella):

I'd like to but i have to turn now. Its pretty late in UK.

OpenStudy (anonymous):

ahh ok